a projectile is fired into the air from the top of a cliff 100m above a valey floor. its initial veloity is 80 m/ at an angle of 25 above the horozontal. assuming free fall conditions dtermine

-the displacment of the projectile after 1 second with respect to the base of the cliff

-total time of flight

-the horozontal distance from the base of the cliff

With the steps please
Answer #1

a)

vertical displacement after 1 s

dy = u sinx * t - 0.5 gt^2

dy = 80 sin 25 - 4.9

d y = 28.91 m

Horizontal displacement

dx = 80 cos 25 * 1 = 72.5 m

net displacement

d^2 = dx^2 + dy^2

d = 78.06 m

======

b)

Considering motion along vertical

h = ho + u sin x t - 0.5 gt^2

0 = 100 + 80 sin 25 * t - 4.9 t^2

solving for t

t = 9.483 s

======

c)

Horizontal range

R = u cos x * t

R = 80 cos 25 * 9.483

R = 687.56 m

=======

Comment in case any doubt, will reply for sure . Goodluck